# High-order WENO finite-difference methods for hyperbolic nonconservative systems of Partial Differential Equations

### v1

@@ 6.1.1 Small perturbations of a steady-state solution
$\left(h_{1},h_{1}u_{1},h_{2},h_{2}u_{2}\right)=\begin{cases}(1+\sigma,0,-1-Z(x),0),&\text{ if }0.1<x<0.2,\\ (1,0,-1-Z(x),0),&\text{ otherwise. }\end{cases}$
The computational domain is $[-0.2,1]$ with extrapolation boundary conditions. We have first checked that for $\sigma=0$ the initial condition is preserved to machine accuracy, i.e. that the C-property is satisfied. Next, we set $\sigma=0.00001$. The numerical solutions computed with $200$ points are compared with a reference solution computed with $2000$ points. The final time is $T=0.15$. The numerical water surface, i.e. $\eta_{1}=-Z+h_{1}+h_{2}$, corresponding to the smooth and the discontinuous topographies are reported in Figure 1 (left and right respectively): it can be seen that small waves are captured without spurious oscillations as expected.
Figure 1: Section 6.1.1: small perturbation of a water-at-rest stationary solution with smooth (left) and discontinuous (right) bottom topographies. The numerical solution was obtained with 200 points and the reference solution was computed with 2000 points: water surface at $T=0.15$.
@@ 6.1.4 Internal dam-break with non-flat bottom
The constant gravitational acceleration is $g=9.81$ and the density ratio is $r=0.998$. The computational domain is $[-5,5]$. The steady state obtained with $500$ points is shown in Figure 5. The converged results agree well with areference solutions computed with $2000$ grids and with those in [22].
@@
6.1.5 Internal dam break with loss of hyperbolicity
This test is taken from [24]. The constant density ratio is $r=0.98$ and the gravitational constant is $g=9.8$. The computational domain is $[0,10]$. The final time is $T=20$. The bottom topography and the initial condition are given now by
$Z(x)=0.47e^{-(x-5)^{2}},$
$h_{2}=\left\{\begin{array}[]{ll}0.5-Z(x),&\text{ if }x<5,\\ 0.97-Z(x),&\text{ otherwise, }\end{array}\quad h_{1}=1-h_{2}-Z(x),\quad u_{1}=u_{2}=0.\right.$
In this test, Methods 1 and 2 give different solutions due to the corrections of hyperbolicity losses: Figure 6 shows the numerical results computed with Methods 1 and 2 using $200$ points. The points in which a hyperbolicity loss has been detected are marked in purple.
Figure 6: Section 6.1.5: internal dam break with loss of hyperbolicity at $T=20$. The contrast between Method 1 and Method 2. Left: Method 1, Right: Method 2.
It can be observed that fewer corrections are needed with Method 1 (4 points) than with Method 2 (7 points). This is so during the whole simulation, what seems to indicate that Method 1 has better properties of hyperbolicity preservation.
@@ 6.1.6 Accuracy test
The computational domain is $[-\pi/2,3\pi/2]$ and periodic boundary conditions are considered. We compute the numerical solutions until $T=0.1$. We use 3rd, 5th, and 7th order WENO with linear weights for this smooth problem to check the numerical accuracy. The reference solutions for 3rd and 5th are computed by each order WENO scheme with 6400 uniform grids respectively. We use 12800 grids to obtain the reference solution for the 7th order accuracy test to get a more accurate reference. The third-order TVD Runge Kutta is used in all cases for time stepping and, in order to reach the same order of accuracy in time and space, the time-step is set to $\Delta t^{\frac{k}{3}}$, where $k$ is the WENO order and $\Delta t$ is the step given by the CFL condition.
Tables 1 and 2 show the $L^{\infty}$ error and the empirical order of convergence: in all cases, the optimal order is reached.
Table 1: Section 6.1.6. $L^{\infty}$ error and empirical order of convergence using WENO3 & WENO7.
N 3rd order test 7th order test
$h_{1}$ error order $h_{1}+h_{2}$ error order $h_{1}$ error order $h_{1}+h_{2}$ error order
25 3.32e-1 5.84e-3 1.62e-1 3.37e-3
50 1.10e-2 1.60 2.10e-3 1.48 1.20e-2 3.76 6.26e-4 2.43
100 2.08e-2 2.40 4.96e-4 2.08 9.58e-4 3.65 1.82e-5 5.10
200 3.19e-3 2.71 7.43e-5 2.74 1.94e-5 5.63 1.82e-7 6.64
400 4.24e-4 2.91 9.75e-6 2.93 2.04e-7 6.57 1.49e-9 6.94
800 5.31e-5 3.00 1.23e-6 2.99 1.64e-9 6.96 1.01e-11 7.20
Table 2: Section 6.1.6. Errors in $L^{\infty}$ norm and empirical order of convergence for WENO5.
N $h_{1}$ error order $h_{1}+h_{2}$ error order
25 5.38e-2 1.42e-3
50 4.65e-3 3.53 4.50e-4 1.66
100 4.84e-4 3.26 1.42e-5 4.98
200 1.95e-5 4.63 5.52e-7 4.69
400 5.92e-7 5.04 1.74e-8 4.99
800 1.76e-8 5.07 4.33e-10 5.33
@@
We have used this test to compare the computational costs of Methods 1 and 2: the CPU times corresponding to a computation with $200\times 200$ points are shown in Table 4. As expected, Method 1 is more expensive, the speedup ratio being of around 1.3. Similar speedups are observed in other test cases.
@@
Table 4: Section 6.2.1. Methods 1 and 2: CPU times and speedup.
@@ 6.2.2 Interface propagation with flat bottom
where $\Omega=\{x<-0.5,y<0\}\cup\left\{(x+0.5)^{2}+(y+0.5)^{2}<0.25\right\}\cup\{x<0,y<-0.5\}$. The constant gravitational acceleration is $g=10$ and the density ratio is $r=0.98$. The flat bottom topography is given by $Z(x,y)\equiv-1$. The computational domain is $[-1,1]\times[-1,1]$ and the final time, $T=0.1$. The computational results obtained with two meshes of $400\times 400$ and $800\times 800$ are shown in Figure 7.
@@ 6.2.3 Interface propagation with non-flat bottom
The computational domain is $[-1,1]\times[-1,1]$ and the values of $g$ and $r$, and the final time are the same. We show again the results computed with $400\times 400$ and $800\times 800$ points in Figure 8. The numerical solutions are in good agreement with those presented in [33] but finer structures of the flow are actually captured here.
@@ 6.2.4 Internal circular dam break over flat bottom topography
In this test, we consider $g=9.81$, $r=0.998$, and $Z\equiv-2$. The final time is $T=20$.
Figure 9: Section 6.2.4: internal circular dam breaking in 2D with a flat bottom. Numerical solutions computed with $200\times 200$ grids. Diagonal slices of the water surface $h_{1}+h_{2}+Z$, interface $h_{2}+Z$ and bottom at times $t=4,6,10,14,16,20$ (from top to down and from left to right).
Diagonal slices $y=x$ of the numerical results computed with $200\times 200$ points for times $t=4,6,10,14,16,20$ are shown in Figure 9. We show the contour lines of the water interface for times $t=4,20$ in Figure 10. It is worth mentioning that the results agree well with those in [9, 22].
Figure 10: Section 6.2.4: internal circular dam breaking in 2D with a flat bottom. Numerical solutions computed with $200\times 200$ grids. Contour lines of the interface $h_{2}+Z$ at times $t=4$(left), $t=20$(right).
@@ 7 Conclusion
A new family of high-order WENO finite-difference methods to solve hyperbolic nonconserative PDE systems have been proposed. These methods are based in a general strategy in which, instead of reconstructing fluxes using a WENO operator, what we reconstruct is the nonconservative products of the system which are computed using the selected family of paths. Moreover, if a Roe linearization is available, the nonconservative products can be computed through matrix-vector operations instead of path-integrals. The main advantages of this theory are the following:
  • the high-order accuracy in space only depends on the order of the selected WENO operator and no extra order-dependent correction terms are required;
  • no integrals in the cells have to be computed so that reconstructions with uniform accuracy in the entire cells are required.
The methods have been extended to systems with source terms and two strategies have been introduced to design well-balanced methods. This methodology has been successfully applied to obtain high-order numerical methods for the 1D and 2D two-layer shallow water equations that preserve water-at-rest steady states. A number of numerical tests confirm the high-order accuracy of the methdos as well as their shock-capturing and well-balanced properties.
The second-strategy introduced here to design well-balanced methods can be combined with the technique developed in [27] to obtain numerical methods that preserve both water-at-rest and moving equilibria for the 1D shallow-water model or some particular families of stationary solutions in the 2D case: this will be done in an upcoming work.
@@ Acknowledgements
The work of B. Ren is supported by the China Scholarship Council. The work of C. Parés is supported by Spanish projects RTI2018-096064-B-C1 and PID2022-137637NB-C21 funded by MCIN/AEI/10.13039/501100011033 and FSE+.

### v2

@@ 6.2.1 Small perturbations of a steady-state solution
$\left(h_{1},h_{1}u_{1},h_{2},h_{2}u_{2}\right)=\begin{cases}(1+\eta,0,-1-Z(x),0),&\text{ if }0.1<x<0.2,\\ (1,0,-1-Z(x),0),&\text{ otherwise. }\end{cases}$
The computational domain is $[-0.2,1]$ with extrapolation boundary conditions. We have first checked that for $\eta=0$ the initial condition is preserved to machine accuracy, i.e. that the C-property is satisfied. Next, we set $\eta=0.00001$. The numerical solutions computed with $200$ points are compared with a reference solution computed with $2000$ points. The final time is $T=0.15$. The numerical water surface, i.e. $\eta_{1}=-Z+h_{1}+h_{2}$, corresponding to the smooth and the discontinuous topographies are reported in Figure 3 (left and right respectively): it can be seen that small waves are captured without spurious oscillations as expected.
Figure 3: Section 6.2.1: small perturbation of a water-at-rest stationary solution with smooth (left) and discontinuous (right) bottom topographies. The numerical solution obtained with 200 points is compared with the reference solution obtained with 2000 points: water surface at $T=0.15$.
@@ 6.2.4 Internal dam-break with non-flat bottom
The constant gravitational acceleration is $g=9.81$ and the density ratio is $r=0.998$. The computational domain is $[-5,5]$. The steady state obtained with $500$ points is shown in Figure 7. The converged results agree well with reference solutions computed with $2000$ grids and with those in [26].
@@ 6.2.5 Accuracy test
The computational domain is $[-\pi/2,3\pi/2]$ and periodic boundary conditions are considered. We compute the numerical solutions until $T=0.1$. We use 5th, and 7th order WENO-Z for this smooth problem to check the numerical accuracy. The reference solutions for 5th are computed by each order WENO scheme with 6400 uniform grids. We use 12800 grids to obtain the reference solution for the 7th order accuracy test to get a more accurate reference. The third-order TVD Runge Kutta is used in all cases for time stepping and, in order to reach the same order of accuracy in time and space, the time-step is set to $\Delta t^{\frac{k}{3}}$, where $k$ is the WENO order and $\Delta t$ is the step given by the CFL condition.
Tables 2 show the $L^{\infty}$ error and the empirical order of convergence: in all cases, the optimal order is reached.
Table 2: Section 6.2.5. $L^{\infty}$ error and convergence rates using WENO3, WENO5, and WENO7
N 3th order test 5th order test 7th order test
$h_{1}$ $h_{1}+h_{2}$ $h_{1}$ $h_{1}+h_{2}$ $h_{1}$ $h_{1}+h_{2}$
25 3.32e-1   — 5.84e-3   — 5.38e-2   — 1.42e-3   — 1.62e-1   — 3.37e-3   —
50 1.10e-2 (1.60) 2.10e-3 (1.48) 4.65e-3 (3.53) 4.50e-4 (1.66) 1.20e-2 (3.76) 6.26e-4 (2.43)
100 2.08e-2 (2.40) 4.96e-4 (2.08) 4.84e-4 (3.26) 1.42e-5 (4.98) 9.58e-4 (3.65) 1.82e-5 (5.10)
200 3.19e-3 (2.71) 7.43e-5 (2.74) 1.95e-5 (4.63) 5.52e-7 (4.69) 1.94e-5 (5.63) 1.82e-7 (6.64)
400 4.24e-4 (2.91) 9.75e-6 (2.93) 5.92e-7 (5.04) 1.74e-8 (4.99) 2.04e-7 (6.57) 1.49e-9 (6.94)
800 5.31e-5 (3.00) 1.23e-6 (2.99) 1.76e-8 (5.07) 4.33e-10 (5.33) 1.64e-9 (6.96) 1.01e-11(7.20)
@@
We have used this test to compare the computational costs of Methods 1 and 2: the CPU times corresponding to a computation with $200\times 200$ points are shown in Table 4. The computational costs of Strategy 1 and 2 are comparable while Strategy 1 is slightly more computationally expensive.
@@
Table 4: Section 6.3.1. Methods 1 and 2: CPU times and speedup
@@ 6.3.2 Interface propagation with flat bottom
where $\Omega=\{x<-0.5,y<0\}\cup\left\{(x+0.5)^{2}+(y+0.5)^{2}<0.25\right\}\cup\{x<0,y<-0.5\}$. The constant gravitational acceleration is $g=10$ and the density ratio is $r=0.98$. The flat bottom topography is given by $Z(x,y)\equiv-1$. The computational domain is $[-0.55,0.7]\times[-0.55,0.7]$ and the final time, $T=0.1$. The computational results obtained with two meshes of $400\times 400$ and $800\times 800$ are shown in Figure 8.
@@ 6.3.3 Interface propagation with non-flat bottom
The computational domain, the values of $g$ and $r$, and the final time are the same. We show again the results computed with $400\times 400$ and $800\times 800$ points in Figure 9. The numerical solutions are in good agreement with those presented in [38] but finer structures of the flow are actually captured here.
@@ 6.3.4 Internal circular dam break over flat bottom topography
In this test, we consider $g=9.81$, $r=0.998$, and $Z\equiv-2$. The final time is $T=20$. We show the contour lines of the water interface for times $t=4,20$ in Figure 10. Diagonal slices $y=x$ of the numerical results computed with $200\times 200$ points for times $t=4,6,10,14,16,20$ are shown in Figure 11. It is worth mentioning that the results agree well with those in [13, 26].
Figure 10: Section 6.3.4: internal circular dam breaking in 2D with a flat bottom. Contour lines of the interface $h_{2}+Z$ at times $t=4$(left), $t=20$(right).
Figure 11: Section 6.3.4: internal circular dam breaking in 2D with a flat bottom. Numerical solutions computed with $200\times 200$ grids. Diagonal slices of the water surface $h_{1}+h_{2}+Z$, interface $h_{2}+Z$ and bottom at times $t=4,6,10,14,16,20$ (from top to bottom and from left to right).
@@ 7 Conclusion
A new family of high-order WENO finite-difference methods to solve hyperbolic nonconserative PDE systems have been proposed. These methods are based on a general strategy in which, instead of reconstructing fluxes using a WENO operator, what we reconstruct is the nonconservative products of the system which are computed using the selected family of paths. Moreover, if a Roe linearization is available, the nonconservative products can be computed through matrix-vector operations instead of path-integrals. The main advantages of this theory are the following:
  • the high-order accuracy in space only depends on the order of the selected WENO operator, provided that the path satisfies the symmetry property;
  • no integrals in the cells have to be computed so that reconstructions with uniform accuracy in the entire cells are required;
  • a unified WENO framework that treats non-conservative equations consistently with conservative equations. The framework is compatible with state reconstruction techniques.
The methods have been extended to systems with source terms to design well-balanced methods. This methodology has been successfully applied to obtain high-order numerical methods for the 1D and 2D two-layer shallow water equations that preserve water-at-rest steady states. A number of numerical tests confirm the high-order accuracy of the methods as well as their shock-capturing and well-balanced properties.
The second strategy introduced here to design well-balanced methods can be combined with the technique developed in [30] to obtain numerical methods that preserve both water-at-rest and moving equilibria for the 1D shallow-water model or some particular families of stationary solutions in the 2D case: this will be done in an upcoming work.
@@
Appendix A Fifth-order WENOZ reconstruction
Let us recall here the expression of the fifth-order WENOZ reconstruction used to compute ${F}_{i+1/2}$ as an example. The expression of ${F}_{i-1/2}$ can be obtained then using the mirror principle. Given $F_{j}=F(U_{j})$, $j=i-2,i-1,i,i+1,i+2$, ${F}_{i+1/2}$ is computed as follows:
${F}_{i+1/2}=\sum_{k=0}^{2}\omega_{k}{F}_{i+1/2}^{k},$ (92)
where
$F_{i+1/2}^{0}=\frac{1}{6}(2F_{i-2}-7F_{i-1}+11F_{i}),F_{i+1/2}^{1}=\frac{1}{6}(-F_{i-1}+5F_{i}+2F_{i+1}),F_{i+1/2}^{2}=\frac{1}{6}(2F_{i}+5F_{i+1}-F_{i+2}),$
are third-order interpolation formulas computed in 3 $substencils$, and $\omega_{k}$, $k=0,1,2$ are nonlinear weights to be computed. In the classical WENO-JS scheme [33] the nonlinear weights are given by
$\omega_{k}^{(JS)}=\frac{\alpha_{k}}{\sum_{j=0}^{2}\alpha_{j}},\quad\alpha_{k}=\frac{d_{k}}{(\beta_{k}+\varepsilon)^{p}},\quad k=0,1,2.$
Here $d_{0}=1/10$, $d_{1}=6/10$, ${d}_{2}=3/10$, are the ideal weights leading to the global fifth-order interpolation formula
$F_{i+1/2}=\frac{1}{60}(2F_{i-2}-13F_{i-1}+47F_{i}+27F_{i+1}-3F_{i+2});$
$\beta_{k}$ are the smoothness indicators
[…]
that are defined so that the weights are close to the ideal ones in smooth regions but, when a discontinuity is detected in the stencil, the contribution of the sub-stencil containing it is close to 0 (non-oscillatory weights). WENO-Z scheme [11] propose the global smooth indicator $\tau_{5}=|\beta_{2}-\beta_{0}|$ to achieve the optimal accuracy at the critical points,
$\omega_{k}^{(Z)}=\frac{\alpha_{k}}{\sum_{l=0}^{2}\alpha_{l}},\quad\alpha_{k}=d_{k}\left(1+\left(\frac{\tau_{5}}{\beta_{k}+\varepsilon}\right)^{p}\right),\quad k=0,1,2,$
$\varepsilon=10^{-12},p=2$ are used in this study. There is a vast literature related to the computation of optimal order smoothness indicators: see for instance [4, 37].
@@ Acknowledgements
The work of B. Ren is supported by the China Scholarship Council. The work of C. Parés is supported by Spanish projects PDC2022-133663-C21 and PID2022-137637NB-C21 funded by MCIN/AEI/10.13039/501100011033 and FSE+.
In memoriam
[…]